OpenStudy (anonymous):

Find an equation for each tangent to the curve y=1/(x-1) that has slope -1.

OpenStudy (anonymous):

take the derivative, set it equal to \(-1\) and solve for \(x\)

OpenStudy (anonymous):

ohhh okay, thanks i'll try that!

OpenStudy (anonymous):

oh yeah. haha. misunderstood it i gues.

OpenStudy (anonymous):

okay so i tried doing that and the answer i got was wrong.

OpenStudy (anonymous):

this is how i tried doing it...

OpenStudy (anonymous):

cancel out the minus sign from both sides

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

-1/(x-1)(x-1) = -1 hence (x-1)(x-1) = 1 you gt x =0,2

OpenStudy (anonymous):

okay i see that, thanks him1618. but now how do i find the equations of the tangents?

OpenStudy (anonymous):

See..you need a slope and a point dont yu?

OpenStudy (anonymous):

\[-\frac{1}{(x-1)^2}=-1\] \[\frac{1}{(x-1)^2}=1\] \[(x-1)^2=1\] \[x-1=1\]or \(x-1=-1\) so \(x=0\) or \(x=2\)

OpenStudy (anonymous):

for any line you need a slope and a point....see if you already have them

OpenStudy (anonymous):

?

OpenStudy (anonymous):

um (0, -1) and (2, 1)

OpenStudy (anonymous):

yes...do you have the corresponding slopes?

OpenStudy (anonymous):

um wouldn't it just be -1 for both?

OpenStudy (anonymous):

true..now find both the lines

OpenStudy (anonymous):

am i using the y-y1=m(x-x1) formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much! i got the right answers :) @ x=0 y=-x-1 @ x=2 y=-x+3